### v2

# Inter-Speaker Relative Cues for Two-Stage Text-Guided Target Speech Extraction

Abstract
This paper investigates the use of relative cues for text-based target speech extraction (TSE). We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions that are often lost in absolute categorical representations for continuous-valued attributes. Building on this analysis, we propose a two-stage TSE framework in which a speech separation model first generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity. Within this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in case of continuous-valued attributes, considering both classification accuracy and TSE performance. Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues; (ii) the proposed two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics; and (iii) several relative cues, including language, loudness, distance, temporal order, speaking duration, random cues, and all cues, can even surpass the performance of an enrollment-audio-based TSE system. Further analysis reveals notable differences in discriminative power across cue types, providing insights into the effectiveness of different relative cues for TSE.
I Introduction
TSE systems leverage auxiliary information—such as speaker embeddings derived from an enrollment speech sample, spatial location, visual cues from lip movements, and textual descriptions of the target speaker’s attributes—to enable the TSE model to selectively attend to and extract the desired voice in a mixture of multiple speakers [57, 54]. Applications such as smart assistants, hearing aids, and meeting transcription benefit from TSE’s ability to isolate the desired voice, enhancing both accuracy and usability by leveraging speaker-specific cues [57, 54, 40, 2].
Recent years have witnessed rapid advances in target speech extraction, driven by progress in deep learning, self-supervised pretraining, and multimodal modeling. Early works in speaker-dependent extraction exploited the idea of conditioning the separation network on speaker identity, giving rise to architectures such as SpeakerBeam [56] and VoiceFilter [50]. These models employ a reference-guided framework, where a short clean utterance of the target speaker provides a speaker embedding that modulates the separation network to enhance the target voice.
Subsequent developments incorporated pretrained speaker encoders, such as D-Vectors[49], X-Vectors[46], and ECAPA-TDNN[8], to obtain robust speaker embeddings, decoupling speaker modeling from the extraction network. Moreover, multi-modal approaches integrating audio and video cues have gained attention[35, 24, 48, 52], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[34]. These methods align visual speech movements or lip information with audio stream[3, 23], improving robustness in visually grounded TSE scenarios.
Despite these advancements, several challenges remain, including robustness to unseen speakers, limited availability of clean target references in real-world scenarios, and mismatch between the enrollment speech and the mixture input regarding recording conditions, speaking style, or acoustic environment. Beyond acoustic/visual-based guidance, recent research has explored text-guided extraction. These models use a text query (prompt) describing the speech attributes of a target speaker as a condition to guide the extraction of the desired speaker’s speech signal from a mixture, enabling extraction when no enrollment utterance of the target speaker is available. Compared to other cues, text offers greater accessibility, enabling users to provide instructions without recording audio or video. It also provides greater flexibility, as it can convey a wide range of target speaker characteristics, including linguistic cues (e.g., speech transcription), paralinguistic cues (e.g., gender), and acoustic cues (e.g., loudness of speech). In addition, text offers stronger privacy by avoiding the transmission or storage of voice data or other media.
Hao et al. were the first to successfully introduce text-based TSE that used a LLM for extracting semantic information from text queries and integrating it into a speech separation network [15]. Subsequently, Huo et al. proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch, gender, accent, and speaking rate [16]. Similarly, Seki et al. developed a TSE model conditioned on text prompts describing paralinguistic and non-linguistic attributes such as gender, pitch, speaking speed, and loudness, demonstrating that it can overcome the limitations of existing audio-based TSE methods [42].
Previous text-based TSE studies typically represent continuous-valued speech attributes, such as speaking rate, using a small number of discrete categories (e.g., slow, normal, fast) without considering inter-speaker relations. This is usually done by first estimating each speaker’s attribute values in a mixture and then quantizing them into predefined categories. Such coarse quantization can obscure meaningful inter-speaker differences, as within-category variation is discarded, reducing discriminative information. Consequently, speaker-independent cues lose finer-grained distinguishing information when the target and interfering speakers have subtle attribute differences but are assigned to the same category.
To overcome this limitation, our prior work introduced inter-speaker relative cues, which explicitly describe differences between the target and interfering speakers [7]. By formulating attributes relatively rather than as speaker-specific categories, these cues preserve fine-grained distinctions that are critical for target speaker extraction. To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
This approach is motivated by human perception: people perceive continuous-valued speech attributes primarily through comparative listening, rather than by measuring precise physical quantities. Instead of relying on exact values, such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute, listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [31, 30]. Conventional TSE systems that rely on absolute, discretized labels fail to capture these relative distinctions. Subtle but perceptually meaningful differences between speakers can be lost when multiple speakers fall into the same category (e.g., both labeled as “fast” speaking rate). By explicitly encoding inter-speaker comparisons, such as “Speaker A has a faster speaking rate than Speaker B”, relative attribute cues preserve these contrasts, resulting in more discriminative and perceptually aligned conditioning for target speaker extraction. Although this study restricts mixtures to two speakers, the relative cue formulation can be naturally extended to mixtures with more speakers. For instance, in mixtures with three or more speakers, cues such as “fastest speaking rate” or “second-fastest speaking rate” can be used to specify the target speaker. For discrete-valued speech attributes, listeners identify the target speaker by comparing the attribute category with that of the interfering speaker, further illustrating the principle of relative cues rather than speaker-independent cues.
Building upon these insights, the main contributions of this work are summarized as follows:
  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker-independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
  • Proposed Two-Stage Text-Guided TSE Method: Similar to conventional audio-based approaches, our previous text-based TSE system follows a conditioning-and-extraction paradigm that is prone to speaker confusion during inference [55, 26, 33, 10], leading to degraded performance. Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speaker-independent speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts. This approach not only mitigates the speaker confusion inherent in single-stage TSE systems but also provides interpretability, allowing insight into the relative effectiveness of different cues through analysis of the classification accuracy for each cue.
  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of relative versus independent cues on overall target speaker classification and the corresponding TSE performance for comparable continuous-valued attributes. This analysis provides an empirical perspective on the advantages of relative cues over independent cues.
The remainder of this paper is organized as follows. Section II introduces the inter-speaker relative cues and presents their theoretical justification. Section III describes the proposed two-stage method. Section IV presents the dataset, training configuration, evaluation setup, experimental results, and corresponding analysis. Finally, Section V concludes the paper.
II Inter-Speaker Relative Cues
(a) Independent Cue
(b) Relative Cue for Direct Difference
(c) Relative Cue for Percentage Difference
Figure 1: Illustration of quantization schemes for continuous-valued attributes. (a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example). Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute. (b) Relative cues for direct differences: attribute values of the target and interfering speakers are compared with direct difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Regions outside the shaded area represent perceptually distinct differences. (c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Because the comparison is based on relative differences, the Similar region gradually expands from the zero point. Regions outside the shaded area represent perceptually distinct relative differences.
The proposed text-guided TSE system takes a two-speaker mixture as input along with a text prompt which consists of cues indicating the target speaker, aiming to extract the target speech signal. In this work, we employ inter-speaker relative cues, which describe the relative differences in speech attributes between the target speaker and the interfering speaker in the mixture. According to the value type of speech attributes, they can be categorized into discrete-valued and continuous-valued types.
Continuous-valued speech attributes correspond to estimations of specific speech or acoustic properties. In this study, these attributes include:
  1. 1.
    mean $\bm{F_{0}}$, computed as the average fundamental frequency over all voiced frames in an utterance;
  2. 2.
    $\bm{F_{0}}$ span, defined as the difference between the maximum and minimum $F_{0}$} within an utterance;
  3. 3.
    speaker age, treated as a continuous variable in years;
  4. 4.
    speaking duration, defined as the total active speech time (in seconds) of a speaker within the mixture;
  5. 5.
    speaking rate, measured as the number of syllables per minute in the speech signal;
  6. 6.
    RMS energy (dB), computed as $20\log_{10}$ of the root-mean-square amplitude of the speech signal;
  7. 7.
    speaker-to-microphone distance, expressed in meters; and
  8. 8.
    appearance time, defined as the onset time of the first occurrence of a speaker’s speech in the mixture.
Discrete-valued speech attributes correspond to categorical properties that take values from a finite set. In this study, these attributes include:
  1. 1.
    language, indicating the spoken language of the utterance;
  2. 2.
    gender, representing the speaker’s perceived acoustical gender category;
  3. 3.
    emotional state, describing the affective category expressed in the speech signal; and
  4. 4.
    transcription, corresponding to the linguistic content of the utterance in textual form.
We derive each inter-speaker relative cue by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture. Each relative cue encodes the difference along one specific attribute only. Independent cues, in contrast to relative cues, are derived solely from the attributes of an individual speaker—either the target or the interfering speaker—without accounting for inter-speaker relationships.
II-A Independent Cues
The key difference between relative and independent cues lies in how continuous-valued attributes are represented. Let $x_{i}^{(\mathrm{tar})}$ and $x_{i}^{(\mathrm{inf})}$ denote the value of a continuous-valued speech attribute $i$ (e.g., speaking rate) for the target and interfering speakers, respectively. Independent cues represent each continuous-valued attribute using a small number of discrete categories. The attribute space is segmented into these categories based on the distribution of attribute values across all individual target and interference speakers in the training set. Each attribute value $x_{i}$ is quantized as
$\tilde{x}_{i}=\mathcal{Q}_{i}(x_{i}),\quad\tilde{x}_{i}\in\{\mathit{A}_{i,1},\mathit{A}_{i,2},\dots,\mathit{A}_{i,k_{i}}\},$ (1)
where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the total number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ denote the corresponding categories (e.g., low, normal, and high for mean $F_{0}$).
The independent cue category for attribute $i$ is given by $\tilde{x}_{i}$. The mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig. 1(a). Due to the coarse nature of the categories, perceptually noticeable differences in the original continuous-valued attribute may be mapped to the same discrete category (as illustrated by the shaded region in Fig. 1(a)), in which case the attributes of two speakers cannot be distinguished from each other.
II-B Relative Cues
In contrast, relative cues categorize continuous-valued attributes into groups (e.g., slower, similar, and faster for speaking rate) by comparing the attribute values of the two speakers. For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the direct difference
$\Delta x_{i}^{(.)}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by differences in their values. In human perception, changes in stimulus magnitude can be detected through difference thresholds, allowing listeners to perceive relative differences in such attributes between competing sources. Consequently, variations in these quantities can be directly perceived and interpreted without requiring an explicit baseline reference [32, 6]. For attributes such as mean $F_{0}$, $F_{0}$ span, speaking rate, and speaking duration, we employ the percentage difference
$\Delta x_{i}^{\%}=\frac{x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}}{\min\!\left(x_{i}^{(\mathrm{tar})},x_{i}^{(\mathrm{inf})}\right)}\times 100\%.$ (3)
These attributes represent global prosodic properties, for which human listeners are more sensitive to proportional than absolute changes [12, 21, 47], making this formulation better aligned with human auditory perception across different baselines.
For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{(.)},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure. This measure is then compared against the predefined threshold $\theta_{i}$ of each attribute, which are shown in Table I.
TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivating Speech Perceptual Studies
Speech Attribute Threshold ($\theta_{i}$) Findings from Motivating Speech Perception Literature
RMS energy (dB) 3 dB Perception of relative loudness differences in speech is mediated by changes in loudness-related acoustic features, such as excitation strength derived from the glottal source [43]. McShefferty et al. [28] report that the average just-noticeable difference (JND) in speech-to-noise ratio is approximately 3 dB.
Speaker-to-microphone distance 0.5 m Human perception of speaker-to-microphone distance relies on relative variations in distance-dependent acoustic cues, particularly changes in the direct-to-reverberant energy ratio [44, 29, 53].
Speaker age 10 years Perception of speaker age is based on relative differences in age-related vocal characteristics, such as vocal tract resonance and speaking rate. Listeners typically judge speakers as “younger” or “older” rather than estimating absolute age [25, 45, 51].
Mean $F_{0}$ 6% Pitch perception is governed by relative changes in F0. Listeners’ sensitivity to differences in F0 changes in speech and found that for many listeners, average JNDs were around 1.5–2 semitones (6-12%) [1].
$F_{0}$ span 25% Perception of pitch range ($F_{0}$ span) depends on relative variation in pitch-related features. Wider pitch ranges are often associated with high arousal or positive affect, while narrower ranges indicate low arousal [18, 5, 41, 19].
Speaking rate 15% Humans perceive differences in speaking rate through relative tempo changes, with a JND of approximately 5% of the base rate [36]. Descriptors such as “much slower” or “faster” reflect this relative perception [13, 37].
Speaking duration 15% Perception of speech duration follows Weber’s law, whereby the JND scales proportionally with the original duration, indicating reliance on relative temporal differences [14].
Appearance time 0.1 s
Some thresholds are motivated by directly adopting suggested JNDs or by adjusting them to larger values based on findings from prior psychoacoustic and speech perception studies. The remaining thresholds, which either lack suggested JNDs or have no related studies, are defined by us [^1] These thresholds are largely chosen in an ad hoc manner, as no perceptual studies have used exactly the same experimental setup..
The relative cue category $r_{i}$ for attribute $i$ is assigned as
$r_{i}=\begin{cases}\mathit{R}_{i,1},&\Delta x_{i}^{(*)}>\theta_{i},\\[2.0pt] \mathit{Similar},&|\Delta x_{i}^{(*)}|\leq\theta_{i},\\[2.0pt] \mathit{R}_{i,2},&\Delta x_{i}^{(*)}<-\theta_{i},\end{cases}$ (4)
where $\mathit{R}_{i,1}$, $\mathit{Similar}$, and $\mathit{R}_{i,2}$ denote relational categories corresponding, for example, to longer, similar, and shorter speaking duration relative to the interfering speaker. This mapping is illustrated in Fig. 1(b) and Fig. 1(c). The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) is presented in Table II. The total number of samples is nearly balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
TABLE II: Sample counts of relative categories for each continuous attribute in the test set based on defined thresholds
Continuous-valued Attribute R1 Similar R2
RMS energy 2547 4992 2461
Speaker-to-microphone distance 1441 7124 1435
Speaker age 198 605 197
Mean $F_{0}$ 4394 1081 4515
$F_{0}$ span 3799 2376 3813
Speaking rate 3331 2235 3465
Speaking duration 3881 2227 3892
Appearance time 3179 3511 3310
II-C Discrete-Valued Attributes
For discrete-valued attributes, independent cues specify the attribute value (category) of the target speaker (e.g., the language spoken) and do not consider whether it is shared with the interfering speaker. In contrast, relative cues indicate whether the target and interfering speakers share the same attribute value (category) or are distinct with respect to that attribute. Let $d_{i}^{(\mathrm{tar})}$ and $d_{i}^{(\mathrm{inf})}$ denote the values of discrete-valued attribute $d_{i}$ for the target and interfering speakers, respectively, with $i$ indexing a particular attribute (e.g., language). Then the independent cue category $a_{i}$ for the target speaker is defined as
$a_{i}=d_{i}^{(\mathrm{tar})},$
while the relative cue category $r_{i}$ for the target speaker is defined as
$r_{i}=\begin{cases}\text{Same},&d_{i}^{(\mathrm{tar})}=d_{i}^{(\mathrm{inf})},\\[2.84526pt] d_{i}^{(\mathrm{tar})},&d_{i}^{(\mathrm{tar})}\neq d_{i}^{(\mathrm{inf})},\end{cases}$
so that when the attribute values (categories) differ, the relative cue retains the original categorical distinction of the target speaker.
II-D Text Prompt Generation
Inter-speaker relative cues are formulated as natural language descriptions (i.e., text prompts) that explicitly capture the differences between the target and interfering speakers, and are used for both training and evaluation of the TSE systems. For each mixture, a target speaker is specified based on the dataset annotations (see Section IV), and a corresponding prompt is generated by instantiating a predefined template with one or more relative cues. Specifically, one prompt is generated for each available individual cue, as well as for combinations of cues, either randomly selected or comprising all available cues (see details below). In the first experimental setting (Section IV-C), cue categories are chosen to ensure clear inter-speaker contrast, thereby providing sufficient discriminability for target speaker extraction. In contrast, the second experimental setting (Section IV-D) adopts a different definition of cue categories, allowing labels such as “same” or “similar” during testing. The template structure used for prompt generation follows the implementation described in the code [^2] https://github.com/daiwangsnr/Inter-Speaker-Relative-Cues-for-Text-Guided-Target-Speech-Extraction/blob/main/prompt_template.py, incorporating action-oriented verbs such as extract, separate, or isolate, along with descriptive phrases that emphasize the distinguishing characteristics of the target speaker. For example, a cue vector such as (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.”
Prompts are generated in three configurations for each mixture:
  • Individual cues: A separate prompt is generated for each available cue.
  • Random cues: If at least three cues are available, a subset is randomly sampled for each mixture, with size $2$ to n1n-1 (where $n>3$ is the total number of available cues), and used to generate a prompt. This configuration enables robust evaluation under partial and randomly combined cues.
  • All cues: If at least two cues are available, a prompt is generated that incorporates all cues, enabling evaluation under fully combined cue conditions.
We also experimented with GPT-4o-mini to automatically rephrase templates for greater linguistic diversity. Although this increased variation, performance slightly decreased; thus, template-based prompts are retained in our experiments. Table III (Section IV) reports the number of mixtures associated with each relative cue in each subset. Some cue names differ from their corresponding speech attributes (e.g., loudness [^3] In this work, “loudness” refers to RMS energy and does not strictly correspond to its conventional perceptual definition vs. RMS energy, age vs. speaker age, pitch vs. mean $F_{0}$, pitch range vs. $F_{0}$ span, distance vs. speaker-to-microphone distance, temporal order vs. appearance time, and emotion vs. emotional state). These differences arise because some cue names are abbreviated, while others are expressed in more perceptually meaningful terms. Since not all mixtures include every cue (i.e., the target speaker is not necessarily characterized by all possible cues), the counts for individual cues do not sum to the total number of mixtures.
Figure 2: Overview of the proposed two-stage text-guided TSE system during inference.
III Proposed Two-Stage Method
In this work, we adopt a two-stage approach for text-guided target speech extraction, comprising: (1) speech separation and (2) target speech classification guided by a text prompt. This design is inspired by prior two-stage target speech extraction framework, which first separate mixture speech signals that perform speech separation followed by speaker verification based on speaker embeddings derived from enrollment utterances [10]. Such a cascaded design [10] has been shown to improve TSE performance compared to single-stage TSE systems. In this two-stage framework, the two components are trained independently, which simplifies the optimization process and allows each module to focus on a more specific objective. This modular strategy leads to a more stable and interpretable architecture than single-stage TSE approaches. An overview of the system during inference is provided in Fig. 2: the input mixture is first separated into individual speech signals, after which a text-guided target selection mechanism identifies and outputs the desired speech signal.
Let the input mixture be denoted as mixmix, which contains two speakers. Given a text prompt $p$ describing one or more relative cues of the target speaker, the goal is to extract the target speech signal from mixmix.
Stage 1: Speech Separation
In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$ as
$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}({mix}),$
trained with utterance-level permutation invariant training (PIT) [20], where $\hat{s}_{1}$ and $\hat{s}_{2}$ denote the estimated speech signals of the individual speakers. PIT resolves the permutation ambiguity by computing the loss for the two possible output-reference assignments and selecting the minimum. This allows the model to learn speaker-specific separation without assuming a fixed output order. The separation stage produces isolated speech signals that serve as input for the subsequent target speech classification stage.
Stage 2: Target Speech Classification via Text Prompt
The second stage identifies which separated signal corresponds to the target speaker using the provided text prompt.
Speaker and Text Embeddings
Each separated signal $\hat{s}_{j}$ (where $j\in\{1,2\}$) is encoded by an audio encoder to produce a fixed-dimensional embedding as
$\tilde{z}_{j}=\mathcal{A}(\hat{s}_{j}).$
The audio encoder consists of a pre-trained self-supervised speech model (wav2vec2-large-xlsr-53444https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
Similarly, the text prompt $p$ is encoded into the same embedding space using a text encoder $\mathcal{T}$ as
$\tilde{z}_{p}=\mathcal{T}(p),$
which includes an LLM-based model (LLaMA 3.2 1B Instruct [^5] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct) followed by a trainable linear projection with ReLU and layer normalization, mapping the LLM output to the speaker embedding space to align textual and speaker representations in a shared, directly comparable latent space.
Similarity Computation and Classification
The similarity of the text prompt and each separated signal is measured via the cosine similarity of their embeddings as
$\text{sim}_{j}=\frac{\tilde{z}_{p}\cdot\tilde{z}_{j}}{\|\tilde{z}_{p}\|\,\|\tilde{z}_{j}\|},\quad j\in\{1,2\}.$
We define the logit as the difference between the two similarity scores as
$\text{logit}=\text{sim}_{1}-\text{sim}_{2},$
which measures how much more similar the separated signal $\hat{s}_{1}$ is to the text prompt compared to $\hat{s}_{2}$. The logit is scaled by a temperature parameter $T$ ($T<1$) and mapped to a probability using the sigmoid function:
$\hat{y}=\sigma(\text{logit}/T),$
where $\hat{y}\in[0,1]$ represents the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the prompt. A temperature $T<1$ sharpens the output distribution, making predictions more confident (closer to 0 or 1). Empirically, we find that $T=0.2$ yields good performance. The predicted target speaker index is then given by
$\text{pred}=\begin{cases}1,&\hat{y}>0.5,\\ 0,&\text{otherwise}.\end{cases}$
During inference, the signal with the higher similarity score is selected as the target speech $\hat{s}_{\text{tar}}$.
Training Label Generation and Loss
To train the classifier, ground-truth labels are generated for each two-speaker mixture by comparing the separated signals with the reference target speech $s_{\mathrm{ref}}$ using scale-invariant signal-to-distortion ratio (SI-SDR) [22] as
[…]
These labels indicate which separated signal corresponds to the target speaker. The classifier is trained by minimizing the binary cross-entropy (BCE) loss between the predicted probability and the generated label:
$\mathcal{L}=\mathrm{BCE}(\hat{y},y),$
encouraging the model to correctly associate the text prompt with the separated speech signal of the target speaker.
IV Evaluation
The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems and investigate the relative effectiveness of different cues within the two-stage framework, and (ii) examine whether relative cues yield better performance than independent cues. The following sections provide an overview of the experimental dataset and training configurations, followed by detailed descriptions of each experiment, including the setup, results, and analysis.
IV-A Dataset
To ensure broad linguistic and acoustic diversity, we construct audio mixtures using the same corpora as in our prior work[7]. Each mixture contains speech from two speakers—one designated as the target and the other as the interference—together with a text prompt describing their relative characteristics.
(1) Extraction of Speech Attributes: To provide ground-truth labels for the relative cue categories in each mixture, a set of speech attributes is extracted for every utterance. These attributes serve as the basis for constructing inter-speaker relative cues (e.g., higher pitch, faster speaking rate). All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or reverberation.
Mean $F_{0}$ and $F_{0}$ span are extracted using the Librosa pYIN library [^6] https://librosa.org/doc/main/generated/librosa.pyin.html. When computing speaking duration, word- and pause-level durations are first obtained. If transcription annotations are available, these durations are extracted from time boundaries provided by the Montreal Forced Aligner [^7] https://montreal-forced-aligner.readthedocs.io/en/latest/; otherwise, they are estimated using the Silero voice activity detection tool [^8] https://github.com/snakers4/silero-vad. Speaking duration is then defined as the sum of word durations and short pauses, excluding pauses longer than 0.6 s, which are treated as silence. Speaking rate is computed as the total number of syllables divided by the speaking duration. Syllable counts are estimated based on vowel sequences: in Chinese, each character is treated as one syllable, whereas in English, French, German, and Spanish, syllables are approximated by counting contiguous vowel groups within words. The RMS energy is computed from reverberant speech, averaged over active speech segments to account for speaking duration. Since the mixtures include room acoustic simulations, the speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone. Language, gender, age, emotional state, and transcription are obtained directly from the original dataset annotations.
(2) Audio Mixtures: Each corpus is divided into non-overlapping training, validation, and test subsets, with disjoint speakers and no shared utterances. We generated 100,000 mixtures for training and 10,000 mixtures each for validation and testing. To create realistic speech mixtures, each mixture is composed of two speech signals, $S_{1}$ and $S_{2}$, from different speakers. Both signals are first processed to remove initial and trailing silence, using forced alignment for transcribed speech or voice activity detection for untranscribed segments, and are subsequently temporally arranged with partial overlaps to mimic conversational dynamics. If either $S_{1}$ or $S_{2}$ is shorter than 3 s, the overlap duration is set to the length of the shorter signal, and zero-padding is applied at the beginning of the shorter signal using a randomly selected start offset uniformly drawn from $0,\text{duration of longer sample}-\text{duration of shorter sample}]$. For longer signals, the overlap duration is set to $(\text{length}(S_{1})+\text{length}(S_{2})-6)$ s, with $S_{2}$ padded using a fixed start offset of $6-\text{length}(S_{2})$ s.
To simulate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library[9], representing different speaker positions. Room dimensions are randomly sampled within 9–11 m (length/width) and 2.6–3.5 m (height), with RT60 values of 0.3–0.6 s. The microphone is positioned at the room center, and speakers are placed such that horizontal speaker–microphone distances range from 0.3 to 1.5 m and heights from 1.6 to 1.9 m. The reverberant $S_{2}$ is fixed while $S_{1}$ is scaled to a signal-to-interference ratio (SIR) uniformly sampled between –6 and 6 dB. For each mixture, one of the reverberant signals is randomly selected as the target speech.
(3) Text Prompts: The details of text prompts generation are described in Section II, Table III summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power. Since the “random” and “all” cues are not available in every mixture to specify the target speaker, the number of samples associated with these cues does not correspond to the total number of mixtures.
TABLE III: Sample counts of each relative cue in the training, validation, and test sets, excluding cues labeled as “Similar” or “Same”
Relative Cue Training Validation Test
Random Cues 99641 9962 9956
All Cues 99960 9996 9997
Language 73845 7192 7157
Transcription 92311 9186 9189
Gender 49114 5759 5125
Emotion 20377 1935 2012
Speaking Rate 67821 6674 6796
Speaking Duration 76498 7880 7773
Pitch 89366 8917 8909
Pitch Range 77019 7715 7612
Distance 27983 2672 2876
Age 3130 348 395
Loudness 50731 4971 5008
Temporal Order 60625 6686 6489
IV-B Training Configurations
For the speech encoder in the classification model of the proposed two-stage framework, all cue types span linguistic, paralinguistic, and acoustic information, making the encoder central to capturing these diverse representations. We adopt the encoder from the pretrained wav2vec2-large-xlsr-53 model, a multilingual self-supervised speech representation model based on the wav2vec 2.0 architecture [4]. Trained on speech from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, it learns universal, language-independent representations directly from raw audio. Compared to the ECAPA-TDNN speaker encoder, which focuses on speaker-specific embeddings, wav2vec2-large-xlsr-53 captures richer linguistic, paralinguistic, and acoustic features, making it well suited for cross-lingual target speech extraction. We use the first five encoder layers, with fine-tuning applied to the fifth layer.
To encode text prompts into embeddings, we use the LLaMA 3.2 1B Instruct model for both single-stage TSE systems and the classification model in the two-stage framework. The LLaMA 3.2 1B instruction model is fine-tuned with LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
To ensure high-quality signal separation, we adopt the TF-Locoformer [39], which demonstrates state-of-the-art performance and robustness under reverberation. Single-stage TSE and two-stage separation models are trained with negative SI-SDR loss, float16 mixed precision, and a batch size of 24 (24 V100 GPUs, batch size 1 per GPU). The classification models also use float16 precision and batch size 24.
All models are optimized using AdamW [27], with learning rates and gradient clipping configured individually for each model. The dual-path transformer (DPT)-based single-stage TSE model (used in the following subsection) was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and 1000-step linear warm-up. The TF-Locoformer-based single-stage TSE was trained with a learning rate of 1e-4 and a maximum norm of 5, whereas the two-stage separation model used a learning rate of 1e-3 with the same maximum norm. The classification model use a learning rate of $1\times 10^{-4}$ with max norm 30. Different learning rates and maximum norm values were adopted to account for the distinct architectures and convergence behaviors of the respective models. Learning rates are halved if validation loss does not improve for three epochs, and training employs early stopping with a maximum of 100 epochs.
TABLE IV: Comparison of single-stage and two-stage TSE systems across different cues
Cue Type Single-Stage Two-Stage
DPT-based TSE TF-Locoformer TSE TF-Locoformer TSE
SI-SDRi PESQ SI-SDRi PESQ Sep ACC SI-SDRi PESQ
Random Cue $8.4$ $1.78$ $13.2$ $2.70$ 99.299.2% $17.1$ $3.39$
All Cue $8.9$ $1.80$ $14.3$ $2.88$ 99.899.8% $17.4$ $3.41$
Language $5.5$ $1.71$ $11.4$ $2.67$ 99.399.3% $17.3$ $3.41$
Transcription $4.3$ $1.67$ $8.7$ $2.47$ 90.490.4% $12.5$ $3.22$
Gender $8.8$ $1.80$ $12.8$ $2.75$ 98.198.1% $16.7$ $3.37$
Emotion $4.9$ $1.56$ $7.8$ $2.44$ 87.287.2% $10.3$ $3.10$
Pitch $7.9$ $1.77$ $12.8$ $2.75$ 94.694.6% $14.8$ $3.30$
Pitch Range $5.3$ $1.68$ $9.0$ $2.47$ 88.788.7% $11.3$ $3.14$
Loudness $8.7$ $1.80$ $13.7$ $2.84$ 97.297.2% $16.0$ $3.35$
Distance $8.3$ $1.78$ $14.6$ $2.89$ 98.498.4% $17.0$ $3.38$
Age 1.0-1.0 $1.38$ 1.9-1.9 $1.71$ 64.364.3% 1.6-1.6 $2.45$
Temporal Order $8.7$ $1.85$ $15.1$ $2.99$ 99.899.8% $18.0$ $3.45$
Speaking Rate $4.0$ $1.68$ $7.9$ $2.44$ 91.391.3% $13.1$ $3.27$
Speaking Duration $7.4$ $1.79$ $13.0$ $2.80$ 95.995.9% $15.6$ $3.34$
IV-C Two-Stage Method versus Single-Stage Baselines
Experimental Setup: We first evaluate the proposed two-stage method against single-stage baselines. In the first stage, the separation model is trained from scratch to obtain separated speech signals. In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, with each prompt corresponding to one of the fourteen relative cue types listed in Table III. The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model, obtained by adapting the TF-Locoformer speech separation model for target speech extraction. Specifically, the model is reconfigured to predict only the target speaker by reducing the number of output speakers in the transposed convolution layer from 2 to 1. Text prompts categorized as “same” or “similar” are excluded from training, validation, and testing, as they provide limited discriminative power.
We also train and evaluate a classification model to test an audio-based TSE approach. Here, the text encoder in the two-stage system is replaced with the audio encoder. For each mixture, a clean utterance from the same target speaker is randomly selected, ensuring that the same utterance is not always used for a given speaker. This setup evaluates the effectiveness of using an enrollment audio cue instead of textual relative cues for target speaker extraction. The clean utterance is normalized to match the mixture length via truncation or zero-padding: signals longer than the target duration are randomly cropped, while shorter signals are zero-padded with randomly distributed offsets. It is then convolved with a RIR randomly drawn from the RIR set described in Section II, producing the reverberant enrollment utterance. To ensure a fair comparison with textual cues, the same number of mixture samples is used for training and validation. Enrollment utterances are distinct from the corresponding mixtures, and evaluation is conducted on mixture sets of the same size as those used for each textual cue, enabling a direct comparison between audio- and text-based TSE approaches.
Results and Analysis: We evaluate target speech extraction using scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [22] and perceptual evaluation of speech quality (PESQ) [38]. For the TF-Locoformer-based two-stage framework, we report classification accuracy on separated speech (Sep ACC) and overall TSE performance (SI-SDRi/PESQ), computed as the mean score between predicted and ground-truth target speech under different relative cue conditions. Results are summarized in Table IV.
We evaluate three prompt types introduced in Section II (Individual, All, and Random) on their corresponding mixtures. The two-stage framework also supports target speaker enrollment, selecting the separated signal with the highest cosine similarity to the enrollment embedding, enabling direct comparison with text-based cues. TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based counterpart, achieving over 3 dB SI-SDRi and 0.7–1.1 PESQ improvements, except for the age cue, which remains challenging. This suggests that age prediction performance is limited by cue discriminability rather than model architecture. Classification performance across cue types reflects their contribution to single-stage baselines. Compared to single-stage TF-Locoformer-based TSE, the two-stage method shows clear gains: over 3 dB SI-SDRi improvement for random, all, language, transcription, gender, and speaking rate cues, and over 2 dB for most others (except age). It also yields an average PESQ improvement exceeding 0.5, demonstrating strong practical potential.
Some relative cues, such as gender, emotion, and pitch, perform competitively compared with enrollment audio. Interestingly, certain cue types—including all cues, random cues, transcription, loudness, distance, temporal order, and speaking duration—can even outperform the target speaker’s enrollment audio (see Table V, Sep ACC: textual cues vs. enrollment audio). This suggests that enrollment speech may contain confounding information, as it could be recorded in a different environment or exhibit speech characteristics differing from the target speech. These findings highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging discriminative cue types present in the mixture.
Figure 3: Training and Validation Loss Comparison for TF-Locoformer-based single-stage TSE and TF-Locoformer-based Separation models over epochs. The loss is measured such that lower (more negative) is better.
We also present the training and validation loss curves for the TF-Locoformer-based single-stage TSE and TF-Locoformer-based separation models in Fig. 3. From the curves, we observe that the TF-Locoformer-based TSE model tends to overfit easily, whereas the TF-Locoformer-based separation model converges more gradually and achieves a better validation loss. In the early training epochs, the separation model also attains a better training loss, despite using a lower learning rate. These observations highlight the advantage of the two-stage approach.
To establish an oracle benchmark, we train separate classification models on reverberant–clean speech pairs (Section II, Audio Mixtures of Dataset) using both relative cues and enrollment utterances. We use identical prompt–sample pairs as in the separated-speech setting. The evaluation results are reported in Table V.
TABLE V: Performance of classification across different cues
Cue Type Textual Cues Enrollment Audio
Clean ACC Sep ACC Clean ACC Sep ACC
Random Cues 99.3% 99.2% 95.7% 95.7%
All Cues 99.8% 99.8% 95.7% 95.7%
Language 99.5% 99.3% 96.6% 96.6%
Transcription 90.8% 90.4% 95.8% 95.9%
Gender 98.0% 98.1% 98.2% 98.1%
Emotion 85.8% 87.2% 88.4% 88.6%
Pitch 95.2% 94.6% 95.8% 95.8%
Pitch Range 88.4% 88.7% 95.9% 96.0%
Loudness 98.8% 97.2% 95.7% 95.7%
Distance 99.0% 98.4% 96.2% 96.4%
Age 64.0% 64.3% 93.4% 93.4%
Temporal Order 100% 99.8% 95.6% 95.6%
Speaking Rate 91.5% 91.3% 95.1% 95.4%
Speaking Duration 96.3% 95.9% 95.7% 95.6%
Average 95.8% 95.6% 95.8% 95.8%
Interestingly, the accuracy for each cue type (relative cues and enrollment audio) closely matches that of the model trained on speech separated by the TF-Locoformer model, demonstrating the high quality of the separated signals. The enrollment audio cue shows slightly more consistent classification performance between clean and separated speech, indicating that speaker identity information is more robust than that conveyed by some relative cues, as expected. The variation in classification accuracy across cue types also provides insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
Figure 4: classification accuracy as a function of the attribute difference between target and interfering speakers for each cue type.
We further analyze classification accuracy as a function of the continuous-valued attribute difference between paired target–interference clean speech samples for each cue type. The temporal order cue is excluded due to its perfect classification accuracy (100%). For the remaining cues, logistic regression is used to obtain a smooth estimate of overall classification accuracy over the continuous attribute space. As shown in Fig. 4, accuracy exhibits a clear increasing trend as attribute differences grow. Except for age, all cues reach near-perfect classification when the differences become sufficiently large.
IV-D Relative Cues versus Independent Cues
Experimental Setup: We then evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework. For discrete-valued attributes, relative and independent cues provide the same level of discriminability; therefore, we restrict our evaluation to originally continuous-valued attributes. In total, eight cues are considered, of which two—age and temporal order—are excluded. The age cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable. Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (young: 7063; youth: 10310; middle: 2627; senior: 0). The temporal order cue, defined as the difference in onset time between the target and interfering speech, is inherently relational and therefore does not admit a meaningful absolute categorization. Consequently, the evaluation focuses on the remaining six cues: pitch, pitch range, loudness, distance, speaking rate, and speaking duration.
Separated speech samples are selected from the TF-Locoformer separation model based on mixtures associated with these cue types. Two binary classification models are trained: one guided by relative-cue text prompts and the other by independent-cue text prompts.
Independent Cues
Independent cue categories are defined based on the distribution of attribute values in the training samples (including both target and interference speech), following previous studies[16, 17]. We use equal-frequency binning, where each attribute is quantized into two or three categories with similar amount of samples per each category:
  • Pitch: low / normal / high
  • Pitch range: narrow / normal / wide
  • Loudness: quiet / normal / loud
  • Distance: near / far
  • Speaking rate: slow / normal / fast
  • Speaking duration: short / long
Relative Cues
Relative cue categories are defined based on the process explained in Section II and the thresholds specified in Table I:
  • Pitch: lower / similar / higher
  • Pitch range: narrower / similar / wider
  • Loudness: quieter / similar / louder
  • Distance: nearer / similar / farther
  • Speaking rate: slower / similar / faster
  • Speaking duration: shorter / similar / longer
TABLE VI: Metrics for Independent and Relative Cues
Independent Cues Relative Cues
Cue Type Group N ACC SI-SDRi PESQ N ACC SI-SDRi PESQ
Overall non-same / non-similar 34501 93.5% $14.3$ $3.29$ 38974 94.1% $14.5$ $3.29$
Overall all 59009 75.3% $4.3$ $2.87$ 59009 79.2% $6.3$ $2.93$
Distance non-same / non-similar 5230 94.0% $14.4$ $3.29$ 2876 98.5% $17.0$ $3.38$
Distance all 10000 72.9% $2.9$ $2.81$ 10000 64.1% 2.1-2.1 $2.53$
Loudness non-same / non-similar 6994 91.8% $13.0$ $3.23$ 5008 97.9% $16.3$ $3.36$
Loudness all 10000 79.2% $6.1$ $2.94$ 10000 74.4% $3.7$ $2.85$
Pitch non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
Pitch all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
Pitch Range non-same / non-similar 6652 89.7% $12.0$ $3.19$ 7612 89.6% $11.9$ $3.17$
Pitch Range all 9988 77.0% $5.0$ $2.90$ 9988 80.0% $6.5$ $2.93$
Speaking Rate non-same / non-similar 5979 90.8% $13.0$ $3.26$ 6796 89.8% $12.3$ $3.23$
Speaking Rate all 9031 75.9% $4.9$ $2.91$ 9031 80.1% $7.0$ $2.99$
Speaking Duration non-same / non-similar 3016 99.5% $18.6$ $3.50$ 7773 96.7% $16.0$ $3.36$
Speaking Duration all 10000 64.2% 1.2-1.2 $2.61$ 10000 86.3% $10.4$ $3.10$
We use audio mixtures from the dataset described in Section IV that are annotated with the six relative cues—pitch, pitch range, loudness, distance, speaking rate, and speaking duration. We reuse the same predefined templates structure as described in Section II to generate prompts for independent cues. To isolate the effect of cue representation, we control for linguistic variability by regenerating relative-cue prompts using the same predefined templates as those used for independent cues. This ensures that both prompt types share identical sentence structure and phrasing, differing only in whether the cue is expressed in relative or independent terms. Consequently, any performance differences can be attributed specifically to the cue formulation rather than to variations in prompt wording or style. For example, a regenerated relative-cue prompt for a given target speech sample is “Please extract the speaker with a higher pitch in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch in the audio.”
During training, pairs of audio samples and prompts in which both speech signals belong to the same absolute category or the “similar” relative category are excluded. During evaluation, all pairs of audio samples and prompts are retained, with target labels derived from the PIT results of the separation model. Classification accuracy is reported under different conditions. For independent cues, three conditions are considered: (i) same category, (ii) adjacent categories (e.g., slow–normal for speaking rate), and (iii) distinctly different categories (e.g., slow–fast for speaking rate). For relative cues, two conditions are considered: (i) same category (“similar”) and (ii) different categories. We further analyze cases in which target–interference pairs fall into different relative-cue categories but are quantized into the same, adjacent, or distinct absolute categories, highlighting scenarios where independent cues provide limited discriminability while relative cues remain informative.
Results and Analysis: We compare overall classification accuracy across cue types to evaluate the effectiveness of relative versus independent cues for TSE. For independent cues, pairs in the same category are labeled “same,” and those in different categories “non-same.” For relative cues, pairs in the same category are labeled “similar,” and those in different categories “non-similar.” Results are reported over six continuous-valued attributes.
For independent cues, we report classification accuracy along with SI-SDRi and PESQ for the “non-same” and “all” groups. For relative cues, we report these metrics for the “non-similar” and “similar” groups. We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues. Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples. The results are summarized in Table VI.
The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues. Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information. In particular, for pitch, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
(a) Independent Speaking Rate Cue
(b) Relative Speaking Rate Cue
Figure 5: Classification accuracy and corresponding sample distribution across regions for independent (a) and relative (b) speaking rate cues, where $x^{(\mathit{tar})}$ (y-axis) and $x^{(\mathit{inf})}$ (x-axis) represent the speaking rate values of the target and interfering speakers, respectively.
For the distance and loudness cues, using relative categories assigns more samples to the “similar” group than absolute categories. This is due to the limited ranges of simulated speaker-to-microphone distances and SIR values. In our dataset, distances range from 0.3 to 1.5 m, with a 0.5 m threshold for distinguishing speakers, leading many pairs to be classified as having similar distances. A similar effect is observed for the loudness cue: loudness is approximated using SIR, which is constrained to the range of -6 to 6 dB in our simulation. With a loudness cue threshold of 3 dB, nearly half of the samples are categorized as having similar loudness.
It is informative to examine the specific categories of target–interference samples. Fig. 5 presents the classification accuracy and sample distribution for both independent (Fig. 5(a)) and relative (Fig. 5(b)) cues, using speaking rate as an example. Compared with the relative speaking rate cue, the independent cue results in more samples being categorized into the “same” group (slow–slow, normal–normal, fast–fast), which cannot be distinguished. In contrast, the relative cue preserves the fine-grained distinctions among these samples, supporting the assumption in Section II. For the independent speaking rate cue, accuracy is highest when target and interference fall into clearly distinct categories (e.g., slow–fast or fast–slow), and decreases for adjacent categories (e.g., slow–normal or normal–fast), as smaller relative differences make classification more difficult.
Figure 6: Comparison between independent and relative cue categorizations for the speaking rate attribute.
(a) Independent Speaking Rate Cue
(b) Relative Speaking Rate Cue
Figure 7: Joint distribution of classification accuracy for independent (a) and relative (b) speaking-rate cues across all paired target–interference samples. The x-axis and y-axis correspond to the speaking rates (syllables per minute) of the interfering and target speakers, respectively. Red points indicate correctly classified samples, while blue points indicate incorrectly classified samples. The gray region denotes target–interference speaking-rate differences within $\pm 15\%$.
To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as “non-similar” and examine whether they fall into the same, adjacent, or distinct absolute categories. Using speaking rate as an example (Fig. 6), we report classification accuracy for these conditions on both the full set and the relative-different subset, along with their sample distribution. Accuracy increases with speaking rate difference: in the same condition, it is 48.9% but improves to 79.4% for “non-similar” pairs, indicating added discriminative power from relative comparison. In the adjacent condition, accuracy rises from 87.3% to 89.5%, while in the distinct condition it reaches 98.3% in both cases. These results show that performance strongly correlates with speaking rate difference, with relative cues particularly benefiting cases where absolute quantization obscures pairwise distinctions.
Fig. 7 further illustrates the joint distribution of target and interference speaking rates under absolute and relative categorizations. Compared with independent cues, relative cues yield more correct classifications in the slow–slow, normal–normal, and fast–fast regions, highlighting their advantage when absolute categories are identical.
V Conclusion
We have presented a comprehensive study on text-guided target speech extraction that leverages inter-speaker relative cues across a wide range of speech attributes. This is achieved through theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative versus independent cues on comparable continuous-valued attributes.
Experimental results demonstrate that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE. Analysis of classification accuracy across cue types indicates that, while some cues have strong discriminative power, certain textual cues are even more effective than conventional audio-based methods, highlighting their practical potential. Furthermore, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction. Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of incorporating advanced separation models into TSE frameworks.
Overall, our findings highlight the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE. The proposed two-stage framework not only mitigates the speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.